Solve each system of equations by using substitution
1. 3x + 7y = 19
x+y=5
2. x+3y = 12
5x+y = 4

Answer by jorel1380(3719)

3x + 7y = 19
x+y=5
x=5-y
3(5-y)+7y=19
15-3y+7y=19
4y=4
y=1
x=4
--------------------
x+3y = 12
5x+y = 4
x=12-3y
5(12-3y)+y=4
60-15y+y=4
14y=56
y=4
x=0..